### v2

@@ Abstract.
ccs: Computer systems organization Distributed architecturesccs: Security and privacy Distributed systems security
@@ 7. The Banyan Algorithm
    if ${\left|\textsc{supp}(b)\right|}+{\left|\textsc{supp}(\textsc{nonLeaderBlocks}(k))\right|}>f+p$, then $b$ is unlocked.
@@ 4. Addition 2
    When a block is proposed, an unlock proof of the parent block and a fast vote for the current block is broadcast alongside the proposed block (32 and 29).
@@ 6. Addition 4
Definition 7.6 (Unlock Proof).
An unlock proof (Addition 1) is the collection of valid fast votes, that prove that $b$ is unlocked according to Definition 7.5.
@@
It is possible to omit sending a corresponding notarization vote when a fast vote is sent. A notarization then consists of two multi-signatures, on for notarization and one for fast votes. For the sake of simplicity, we do not consider this version of Banyan in our description and analysis.
@@
21:     $D\leftarrow\emptyset$; $\triangleright$ Set of disqualified ranks
22:     $N\leftarrow\emptyset$; $\triangleright$ Set of blocks for which a notarization vote was sent
23:
24:upon $\neg proposed\;\textbf{and}\;clock()\geq t_{0}+\Delta_{prop}(r_{u})$ do
25:     create a new round k block:
26:     $b\leftarrow(k,u,hash(b_{p}),payload,{signature}_{u})$
27:     $proposed\leftarrow\text{\tt True}$
28:     if $r_{u}=0$ then
29:         broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s unlock proof, fast vote for $b$
30:         $fastVoteSent\leftarrow\text{\tt True}$
31:     end if
32:     broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s unlock proof
33:
34:upon exists a valid round $k$ block $b$ of rank $r$ such that $b\notin N,\;\textbf{and}\;r\notin D\;\textbf{and}\;clock\geq t_{0}+\Delta_{notary}(r)\;\textbf{and}\;\nexists\textit{ valid round $k$ block $b^{\prime}$ of rank $r^{\prime}$}\in[r]\setminus D$ do
35:     if $r_{u}\neq r$ then
36:         broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s unlock proof
37:     end if
38:     if some block in $N$ has rank $r$ then:
39:         $D\leftarrow D\cup\{r\}$;
40:     else
41:         $N\leftarrow N\cup\{b\}$
42:         if $\neg fastVoteSent$ then:
43:              broadcast fast and notarization vote for $b$;
44:              $fastVoteSent\leftarrow\text{\tt True}$;
45:         else
46:              broadcast notarization vote for $b$
47:         end if
48:     end if
49:
@@ 8.1. Deadlock Freeness
${\left|\textsc{supp}(\textsc{max}(k))\right|}+{\left|\textsc{supp}(\textsc{nonLeaderBlocks}(k))\right|}$
$={\left|\textsc{supp}(\textsc{blocks}(k))\right|}>n-f>f+p$
@@
Instead, if $\mathcal{S}\neq\emptyset$, there are at least two rank-0 blocks. By Addition 2 (29), they each contain a fast vote from the (Byzantine) leader. Since $u$ also receives fast votes from $n-f\geq 2f+2p-1$ honest replicas, at least $12f+2p+$ 1 fast votes will be received, i.e., ${\left|\textsc{supp}(\textsc{max}(k)\cup\textsc{nonMaxBlocks}(k))\right|}=2f+2p+1$. By the pigeonhole principle, either ${\left|\textsc{supp}(\textsc{max}(k))\right|}\geq f+p+1$ or ${\left|\textsc{supp}(\textsc{nonMaxBlocks}(k))\right|}\geq f+p+1$. In both cases at least one block is unlocked according to Definition 7.5, thus leading to a contradiction. ∎
@@
Each honest replica will receive at least one round $k+1$ block, together with a notarization and unlock proof for the extended parent block (Addition 2, 32). Thus, each honest replica will send a fast vote for one block (Addition 3, 43). Thus, by Lemma 8.1, at least one block round $k+1$ block will be unlocked. As no honest replica moves to a higher round without an unlocked block being notarized (Restriction 2, 4), and replicas keep notarizing blocks, one unlocked block will be notarized. Finally, honest replicas moving from one round to the next broadcast the notarization and unlock proof (Addition 1, 6), guaranteeing that all honest replicas can add a notarized and unlocked block to their block-tree. ∎
@@ 8.2. Safety
${\left|\textsc{supp}(b^{\prime})\right|}+{\left|\textsc{supp}(\textsc{nonLeaderBlocks}(k))\right|}>f+p$
This implies that blocks different from $b$ have received more than $f+p$ fast votes, more than $p$ of which coming from honest replicas. By 42, 43 and 44, each honest replica broadcasts at most one fast vote per round. Thus $b$ received less than $n-p$ fast votes, a contradiction, as $n-p$ are required for FP-finalization (12).
@@
Consider the case where $b$ and $b^{\prime}$ were both explicitly finalized. SP-and FP-finalization imply the reception of ${\lceil\frac{n+f+1}{2}\rceil}$ unique notarization and fast votes for a block, respectively. This implies the existence of two respective Byzantine quorums, a contradiction (Cachin et al., 2011). Thus, without loss of generality, assume $b^{\prime}$ was implicitly finalized. If $b$ was SP-finalized, by Lemma 8.3 $b^{\prime}$ cannot have been notarized, and thus no honest replica would have extended $b^{\prime}$. Instead, if $b$ was FP-finalized, by Lemma 8.4 then $b^{\prime}$ was not unlocked. Thus, honest replicas would have sent fast, notarization, or finalization votes for a block that was not unlocked, a contradiction (see 34 and the validity definition on 18). ∎
@@ 8.3. Liveness
Note that $\Delta_{prop}(0)=0$. Thus, replica $l$ broadcasts $b$ by time $T+\delta$, and other replicas will receive it by $T+2\delta$. By assumption, no other honest replica enters round $k$ before time $T$ and as $\Delta_{notary}(1)\geq 2\delta$, every honest replica will broadcast a fast vote and a notarization vote for $b$ (Addition 3, 43). Therefore, all honest replicas eventually receive at least $n-f$ fast votes for $b$. When $p=f$, these fast votes are enough to FP-finalize $b$ (Addition 4, 12). When $p<f$, ${\lceil\frac{n+f+1}{2}\rceil}\leq n-f$ notarization votes are sufficient to notarize $b$ instead. As argued previously, honest replicas will not have notarized another block and thus will broadcast a finalization vote (8). Eventually, after an additional communication round, honest replicas will receive ${\lceil\frac{n+f+1}{2}\rceil}\leq n-f$ finalization votes, which leads to SP-finalization (12). ∎
@@ 9.2. Methodology
Figure 5. The red triangles mark the coordinates of the 4 AWS datacenters used in Section 9.3, while the yellow crosses show the coordinates of 4 datacenters in the US used in Section 9.4. The 19 different coordinates of the deployed EC2 instances from Section 9.5 are marked by black circles.
@@
As the goal of our evaluation is to measure the precise impact of our changes, we define latency as the average proposal finalization time, measured at the respective proposer using their system clocks. In turn, throughput is measured as the average number of committed bytes per second at any (non-faulty) replica. We measure the protocols’ behavior under varying load by controlling the size of the payload, where the payload is a bit vector randomly generated by the leader.
@@ 9.4. Effect of Crash-faults
In this experiment, we show the effect that crashes have on Banyan and Internet Computer Consensus. We distribute 19 replicas across four US AWS datacenters, as shown by the yellow crosses in Figure 5. Results are presented in Figure 6(d). As has been noted in literature, rotating leader protocols are sensitive to crashes (or malicious behavior), as at least one full timeout duration must expire before progress can be made in the case of a faulty leader. The timeout parameter will crucially impact the results of such scenarios. In the presented experiment, the timeout was set to 3 seconds. As seen in Figure 6(d), there are no penalties in trying to take the fast path. When they are failures, the performance of Banyan is exactly the one of ICC.
@@ 9.5. Global Network
In this experiment, we simulate a worldwide deployment. We include almost all AWS datacenter locations available to us and distribute 19 replicas across the globe, as shown by the black circles in Figure 5. In this setting and with 1MB payloads, the average proposal finalization time measured for the ICC implementation was 384ms. By running Banyan with $f=6,p=1$, the average proposal finalization time is reduced by 5.8% to 362ms for “free”. Further, for Banyan with $f=4,p=4$, the number drops further by 16% to 324ms (Figure 6(e)).

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 24th International Middleware Conference; December 2–6, 2024; Hong Kong, Hong Kongbooktitle: 24th International Middleware Conference (MIDDLEWARE ’24), December 2–6, 2024, Hong Kong, Hong Kongdoi: 10.1145/3652892.3700788isbn: 979-8-4007-0623-3/24/12ccs: Computer systems organization Distributed architecturesccs: Security and privacy Distributed systems security
@@ 7. The Banyan Algorithm
    if ${\left|\textsc{supp}(b)\cup\textsc{supp}(\textsc{nonLeaderBlocks}(k))\right|}>f+p$, then $b$ is unlocked.
@@
Definition 7.6 (Unlock Proof).
An unlock proof is the collection of valid fast votes, that prove that $b$ is unlocked according to Definition 7.5.
@@ 4. Addition 2
    When a block is proposed, an unlock proof of the parent block and a fast vote for the current block is broadcast alongside the proposed block (31 and 28).
@@ 6. Addition 4
It is possible to omit sending a corresponding notarization vote when a fast vote is sent. A notarization then consists of two multi-signatures, one for notarization and one for fast votes. For the sake of simplicity, we do not consider this version of Banyan in our description and analysis.
@@
21:     $N\leftarrow\emptyset$; $\triangleright$ Set of blocks for which a notarization vote was sent
22:
23:upon $\neg proposed\;\textbf{and}\;clock()\geq t_{0}+\Delta_{prop}(r_{u})$ do
24:     create a new round k block:
25:     $b\leftarrow(k,u,hash(b_{p}),payload,{signature}_{u})$
26:     $proposed\leftarrow\text{\tt True}$
27:     if $r_{u}=0$ then
28:         broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s unlock proof, fast vote for $b$
29:         $fastVoteSent\leftarrow\text{\tt True}$
30:     end if
31:     broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s unlock proof
32:
33:upon exists a valid round $k$ block $b$ of rank $r$ such that $b\notin N,\;\textbf{and}\;clock\geq t_{0}+\Delta_{notary}(r)\;\textbf{and}\;\nexists\textit{ valid round $k$ block $b^{\prime}$ of rank $r^{\prime}$, $r^{\prime}<r$}$ do
34:     if $r_{u}\neq r$ then
35:         broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s unlock proof
36:     end if
37:     $N\leftarrow N\cup\{b\}$
38:     if $\neg fastVoteSent$ then:
39:         broadcast fast and notarization vote for $b$;
40:         $fastVoteSent\leftarrow\text{\tt True}$;
41:     else
42:         broadcast notarization vote for $b$
43:     end if
44:
@@ 8.1. Deadlock Freeness
${\left|\textsc{supp}(\textsc{max}(k))\cup\textsc{supp}(\textsc{nonLeaderBlocks}(k))\right|}$
$={\left|\textsc{supp}(\textsc{blocks}(k))\right|}\geq n-f>f+p$
@@
Instead, if $\mathcal{S}\neq\emptyset$, there are at least two rank-0 blocks. By Addition 2 (28), they each contain a fast vote from the (Byzantine) leader. Since $u$ also receives fast votes from $n-f\geq 2f+2p-1$ honest replicas, at least $12f+2p+$ 1 fast votes will be received. This implies that
${\left|\textsc{supp}(\textsc{max}(k))\right|}+{\left|\textsc{supp}(\textsc{nonMaxBlocks}(k))\right|}\geq 2f+2p+1$
By the pigeonhole principle, either ${\left|\textsc{supp}(\textsc{max}(k))\right|}>f+p$ or ${\left|\textsc{supp}(\textsc{nonMaxBlocks}(k))\right|}>f+p$. In both cases at least one block is unlocked according to Definition 7.5, leading to a contradiction. ∎
@@
Each honest replica will receive at least one round $k+1$ block, together with a notarization and unlock proof for the extended parent block (Addition 2, 31). Thus, each honest replica will send a fast vote for one block (Addition 3, 39). Thus, by Lemma 8.1, at least one block round $k+1$ block will be unlocked. As no honest replica moves to a higher round without an unlocked block being notarized (Restriction 2, 4), and replicas keep notarizing blocks, one unlocked block will be notarized. Finally, honest replicas moving from one round to the next broadcast the notarization and unlock proof (Addition 1, 6), guaranteeing that all honest replicas can add a notarized and unlocked block to their block-tree. ∎
@@ 8.2. Safety
${\left|\textsc{supp}(b^{\prime})\cup\textsc{supp}(\textsc{nonLeaderBlocks}(k))\right|}>f+p$
This implies that blocks different from $b$ have received more than $f+p$ fast votes, more than $p$ of which coming from honest replicas. By 38, 39 and 40, each honest replica broadcasts at most one fast vote per round. Thus $b$ received less than $n-p$ fast votes, a contradiction, as $n-p$ fast votes are required for FP-finalization (12).
@@
Consider the case where $b$ and $b^{\prime}$ were both explicitly finalized. SP-and FP-finalization imply the reception of ${\lceil\frac{n+f+1}{2}\rceil}$ unique notarization and fast votes for a block, respectively. This implies the existence of two respective Byzantine quorums, a contradiction (Cachin et al., 2011). Thus, without loss of generality, assume $b^{\prime}$ was implicitly finalized. If $b$ was SP-finalized, by Lemma 8.3 $b^{\prime}$ cannot have been notarized, and thus no honest replica would have extended $b^{\prime}$. Instead, if $b$ was FP-finalized, by Lemma 8.4 then $b^{\prime}$ was not unlocked. Thus, honest replicas would have sent fast, notarization, or finalization votes for a block that was not unlocked, a contradiction (see 33 and the validity definition on 18). ∎
@@ 8.3. Liveness
Note that $\Delta_{prop}(0)=0$. Thus, replica $l$ broadcasts $b$ by time $T+\delta$, and other replicas will receive it by $T+2\delta$. By assumption, no other honest replica enters round $k$ before time $T$ and as $\Delta_{notary}(1)\geq 2\delta$, every honest replica will broadcast a fast vote and a notarization vote for $b$ (Addition 3, 39). Therefore, all honest replicas eventually receive at least $n-f$ fast votes for $b$. When $p=f$, these fast votes are enough to FP-finalize $b$ (Addition 4, 12). When $p<f$, ${\lceil\frac{n+f+1}{2}\rceil}\leq n-f$ notarization votes are sufficient to notarize $b$ instead. As argued previously, honest replicas will not have notarized another block and thus will broadcast a finalization vote (8). Eventually, after an additional communication round, honest replicas will receive ${\lceil\frac{n+f+1}{2}\rceil}\leq n-f$ finalization votes, which leads to SP-finalization (12). ∎
@@ 9.2. Methodology
Figure 5. Locations of the AWS datacenters used in our testbed. Red triangles represent the 4 datacenters employed in Section 9.3, yellow crosses indicate the 4 US datacenters used in Section 9.4, and black dots denote the 19 worldwide datacenters used in Section 9.5.
@@
As the goal of our evaluation is to measure the precise impact of our changes, we define latency as the average proposal finalization time, measured at the respective proposer using their system clocks. Throughput is calculated as the average number of committed bytes per second at any (non-faulty) replica. We measure the protocols’ behavior under varying load by controlling the size of the payload, where the payload is a bit vector randomly generated by the leader.
@@ 9.4. Effect of Crash-faults
In this experiment, we show the effect that crashes have on Banyan and Internet Computer Consensus. We distribute 19 replicas across four US datacenters, as shown by the yellow crosses in Figure 5. Results are presented in Figure 6(d). As has been noted in literature, rotating leader protocols are sensitive to crashes (or malicious behavior), as at least one full timeout duration must expire before progress can be made in the case of a faulty leader. The timeout parameter will crucially impact the results of such scenarios. In the presented experiment, the timeout was set to 3 seconds. As seen in Figure 6(d), there are no penalties in trying to take the fast path. When there are failures, the performance of Banyan is exactly the one of ICC.
@@ 9.5. Global Network
In this experiment, we simulate a worldwide deployment. We include almost all AWS datacenter locations available to us and distribute 19 replicas across the globe, as shown by the black dots in Figure 5. In this setting and with 1MB payloads, the average proposal finalization time measured for the ICC implementation was 384ms. By running Banyan with $f=6,p=1$, the average proposal finalization time is reduced by 5.8% to 362ms for “free”. Further, for Banyan with $f=4,p=4$, the number drops by 16% to 324ms (Figure 6(e)).